OpenStudy (shaik0124):

A biased die has a probability of 1/4 of showing a 5, while the probability of any of 1, 2, 3, 4, or 6 turning up is the same . If three such dice are rolled, what is the probability of getting a sum of atleast 14 without getting a 6 on any die ? Option 1 : 5/24 Option 2 : 9/160 Option 3 : 1/30 Option 4 : 7/160

OpenStudy (dan815):

ok what have u done so far

OpenStudy (mimi_x3):

Umm dan can i ask you a question

OpenStudy (dan815):

to start off you need to see the prob of getting 14, and the prob of getting 14 with 6s

OpenStudy (mimi_x3):

Like rolling a 5 has a probability of 1/4 but is it right to assume that rolling a 6 has a probability of 3/20

OpenStudy (dan815):

yes

OpenStudy (dan815):

ok that method from what i read so far is fine but

OpenStudy (shaik0124):

My work: since the sum should be atleast 14 that means sum could be 14 or 15 or 16 or 17 or 16(6+6+6) since 6 should not be there on any die we are left with 14 or 15 for 14 and 15 the sum of three dice could be 5+5+4 or 5+5+5 now the probability of getting a 5 is 1/4 so the probability of getting 14 is 1/4 *1/4*1/5 and that of 15 is 1/4 *1/4*1/4 total probability is 1/4 *1/4*1/5+ 1/4 *1/4*1/4

OpenStudy (dan815):

u cannot do that in the case where you have more dices, checking the addition cases is not so easy

OpenStudy (shaik0124):

what's the probability of getting a 4??

OpenStudy (dan815):

3/4*1/5

OpenStudy (shaik0124):

there is no other option except adding the sum of 3 dices to get 14 or 15

OpenStudy (shaik0124):

if this is not the method how will u get 14 or 15 in Question it is asked to add the sum of 3 dices

OpenStudy (shaik0124):

i think what i did is right just say may whats the probability of getting a 4 from 1,2,3,4,5,6 should 6 be included to calculate the probability of 4 (in question it is said that 6 should not be used to calculate the sum of 14)

OpenStudy (dan815):

ok just fix prob of 4

ganeshie8 (ganeshie8):

that looks fine with numbers 1-5, you're left with sums of 14 or 15

ganeshie8 (ganeshie8):

probability of getting a 4 has nothing to do with the later problem

ganeshie8 (ganeshie8):

P(5) = 1/4 P(any other number) = 3/20

ganeshie8 (ganeshie8):

the probabbilites are fixed ^ the die wont know your question in advance to change probabilities accordingly

OpenStudy (shaik0124):

how is it 3/20 can u plz expalin

OpenStudy (shaik0124):

@ganeshie8

OpenStudy (dan815):

7/160?

OpenStudy (shaik0124):

@FateRocks8D

OpenStudy (mimi_x3):

The probability of landing on any number from 1-6 is mutually exclusive on a single roll Meaning you can only spin on 1 ...you cant spin on 1 and 2 simultaneously Anyways so we know that probability of landing on every single number from 1 to 6 is 1 cus when you roll a die then you will automatically land on a number from 1-6 so therefore we can conclude that P(1)+P(2)+P(3)+P(4)+P(5)+P(6)=1 When we addd the probabilities of every event it obviously equals 1 Now we know that the probability of landing on a 5 is 1/4 So we can substitute P(5)=1/4 So our equations looks like the following P(1)+P(2)+P(3)+P(4)+(1/4)+P(6)=1 Subtract a 1/4 from both sides we get P(1)+P(2)+P(3)+P(4)+P(6)=3/4 So we know that the rest all have equal likelihood of landing on its number meaning teh liklihood of landing on 1,2,3,4, and 6 are all the same hence P(1)=P(2)=P(3)=P(4)=P(6)=X So lets represent the probability of landing on any of these 5 numbers as X So lets replace P(1), P(2) and so forth with X since they all have the same probability P(1)+P(2)+P(3)+P(4)+P(6)=3/4 X+X+X+X+X=3/4 5X=3/4 \(\frac{5x}{5}=\frac{3}{4} \div 5\) \(x=\frac{3}{4}*\frac{1}{5}=\frac{3}{20}\)

OpenStudy (shaik0124):

yeah answer is 7/160

OpenStudy (mimi_x3):

So the probability of landing on 1,2,3,4 and 5 is 3/20

OpenStudy (mimi_x3):

how did you get that answer?

OpenStudy (dan815):